# Stability of $f(Q,B)$ gravity via Dynamical System Approach: a Comprehensive Bayesian Statistical Analysis

### v4

@@ IV Dynamical System Analysis
$\frac{dY}{dN}$ $=$ $2-3V+\frac{2Z}{3}-\frac{2Zf_{Q}}{3}+3X$ (49b)
−$-2f_{Q}+W-\frac{2f_{Q}^{\prime}}{3}+XZ-YZ\,,$
@@
$\frac{dV}{dN}$ $=$ −$-4V-2ZV\,,$ (49d)
where $\lambda=\frac{\ddot{H}}{H^{3}}$, we will concentrate on the scenario where $f(Q,B)=f_{0}Q^{m}B^{n}$. The model for this scenario can be expressed using the dynamical variables.
@@
and we get the following dependency relation
$W$ $=$ −$-\frac{X}{n}(Z+3)\,,$ (51)
$\Omega_{\text{m}}$ $=$ −$-V-2f_{Q}+W+3X+XZ-Y\,,$ (52)
@@
It is possible to eliminate the equations for $W$, $\Omega_{\text{m}}$ and $\Omega_{\text{DE}}$ from our autonomous system using the relations mentioned and constraint (46), resulting in a set of only four equations
@@
$\frac{dV}{dN}$ $=$ −$-4V-2ZV\,.$ (55d)
It is important to note that a dynamical system has a critical point and this point must be taken into account when analyzing the system:
@@
for existence condition $m=1-n$. We will now analyze the range of value for $n$, which will result in a stable critical point. While we will not explicitly mention the area of instability, whether it is saddle-like or repulsor-like, it is important to note that the critical point is a de Sitter acceleration phase, and therefore, any kind of instability of the critical point is not supported by observations.
@@
$\left\{0,-4,\frac{-n-\sqrt{-2n^{2}-3n}}{n},\frac{-n+\sqrt{-2n^{2}-3n}}{n}\right\}$ (57)
The phase portrait in Fig. 6 shows the behavior of a dynamical system near a stable critical point, $P_{\star}$. As time progresses, trajectories in the phase space tend to move towards $P_{\star}$, indicating that it is an attractor for the system. This convergence from various initial conditions signifies that small perturbations decay over time, returning the system to the stable state at $P_{\star}$. The stability of $P_{\star}$ can be analyzed using the Jacobian matrix evaluated at $P_{\star}$. Stability of $P_{\star}$ is ensured when all eigenvalues of the Jacobian possess negative real parts, which aligns with the observed behavior in the phase portrait. The overall dynamics of the system are governed by the differential equations defining it, and the phase portrait provides a graphical representation to visualize these dynamics and understand the long-term behavior of the system. The zero eigenvalues suggest the presence of a center manifold, which requires further analysis to determine overall stability.
IV.1 Stability analysis for $P_{\star}$ by using Center Manifold Theory (CMT):
Central Manifold Theory (CMT) is a crucial aspect of dynamical systems theory, focusing on the behavior of systems near fixed points. As detailed by Perko [63], CMT provides a framework for understanding the stability of these points, especially when traditional linear stability theory falls short due to the presence of zero eigenvalues. This theory reduces the dimensionality of the system near critical points, allowing for a more manageable analysis of stability. When a system approaches a critical point, it tends to behave according to an invariant local center manifold, denoted as $W^{c}$. This manifold is associated with eigenvalues that have zero real parts, capturing the essential dynamics of the system near equilibrium.
[…]
@@ IV.1 Stability analysis for $P_{\star}$ by using Center Manifold Theory (CMT):
$\dot{y}=By+G(x,y)$
Here, $A$ is a square matrix with $c$ eigenvalues having zero real parts, and $B$ is a square matrix with $s$ eigenvalues having negative real parts and $(x,y)\in R^{c}\times R^{s}$. The functions $F$ and $G$ satisfy $F(0)=G(0)=0$ and their derivatives at zero are also zero. Furthermore, there is a small positive value $\epsilon>0$ and a function $g(x)$ in $C^{r}(N_{\epsilon}(0))$, which defines the local center manifold and satisfies certain conditions:
@@
for $|x|<\epsilon$. The center manifold can be derived using the system of differential equations:
@@
[…]
At the critical point $P_{\star}$, the Jacobian matrix for the autonomous system represented by Eq. (55) is as follows:
$J\left(P_{\star}\right)=\left(\begin{array}[]{cccc}0&1&0&0\\ -3+\frac{3}{n}&-2&-\frac{4}{3}&-3\\ 0&\frac{9}{2n}&0&0\\ 0&0&0&-4\\ \end{array}\right)$ (61)
The eigenvalues of the Jacobian matrix, as presented in (57), are $\lambda_{1}=0$, $\lambda_{2}=-4$, $\lambda_{3}=\frac{-n-\sqrt{-2n^{2}-3n}}{n}$, and $\lambda_{4}=\frac{-n+\sqrt{-2n^{2}-3n}}{n}$. The corresponding eigenvectors are: $[\frac{4n}{9-9n},0,1,0]^{T}$, [−$[-\frac{3n}{11n+3},\frac{12n}{11n+3},-\frac{27}{22n+6},1]^{T}$, $[\frac{2n}{9},\frac{2}{9}\left(\sqrt{-n(2n+3)}-n\right),1,0]^{T}$ and $[\frac{2n}{9},\frac{1}{9}(-2)\left(n+\sqrt{-n(2n+3)}\right),1,0]^{T}$.
Using the CMT, we examine the stability of the critical point $P_{\star}$. By applying the transformation $X=x-\frac{2n}{3(n-1)}$, $Y=y$, $Z=z$, and $V=v$, we shift this critical point to the origin. The resulting equations in the new coordinate system are as follows:
@@
Upon examining the diagonal matrix in relation to the standard form (58), it is clear that the variables $y$, $z$ and $v$ remain stable, while $x$ acts as the central variable. At this critical point, matrices $A$ and $B$ take on the following form:
@@
In the context of CMT, the manifold is characterized by a continuous differential function. Assuming specific functions for the stable variables $y=g_{1}(x)$, $z=g_{2}(x)$, and $v=g_{3}(x)$, we derived the zeroth approximation of the manifold functions using Eq. (59)
@@
where $\mathcal{O}^{2}$ term encompasses all terms that are proportional to the square or higher powers. The following expression gives the center manifold in this scenario
@@
According to the CMT, the critical point $P_{\star}$ exhibits stable behavior for $(n<0)\vee(n>1)$.
Figure 6: This graph shows the behavior of a four-dimensional system (55) simplified to two dimensions. The parameters are set as $Z=0$, $V=0$, and $n=3.012$.
Fig. 6 reveals that a specific point in this two-dimensional representation (the critical point $P_{\star}$) attracts other points over time, suggesting its stable and attractive nature within the full four-dimensional system. Fig. 6 shows the fascinating world of sink trajectories within a dynamical system, visualized through a phase portrait. This point signifies a location where trajectories tend to sink or converge. The critical point $P_{\star}$ is non-hyperbolic due to the presence of zero eigenvalues and can describe the acceleration of the Universe. It is also an attractor solution, stable in the regions $(n<0)\vee(n>1)$ as determined by CMT. The density parameters for radiation, matter, and DE are $\Omega_{\text{r}}=0$, $\Omega_{\text{m}}=2-\frac{4}{n-1}$, and $\Omega_{\text{DE}}=-1+\frac{4}{n-1}$, respectively, satisfying the constraint in Eq. (46). This scenario corresponds to a deceleration parameter $q=-1$ and a total EoS $\omega_{\text{tot}}=-1-\frac{2\dot{H}}{3H^{2}}=-1$, indicating a de Sitter phase and, consequently, an accelerating expansion of the Universe.
@@
Fig. 7 shows the evolution of density parameters for DE, matter, and radiation as a function of redshift. The model parameters $m=-2.011$ and $n=3.012$, which are obtained from the parametrization method using MCMC analysis for CC + Pantheon+ + BAO datasets, are used.
@@
    The radiation density parameter (ΊrsubscriptΊ$_{\text{r}}$) is represented by the cyan line, which remains almost constant at zero for small redshift values. This emphasizes its minimal contribution in the present epoch.
@@
In Fig. 7, we present the evolution of the density parameters for dark energy (DE), matter, and radiation as a function of redshift ($z$), utilizing a model parameter value of $m=-2.011$ and $n=3.012$ obtained from our MCMC analysis. The initial conditions for this plot are $X=10^{11}$, $Y=3.29\times 10^{14}$, $Z=0.008$, and $V=4.54\times 10^{-5}$. The magenta line representing DE exhibits a significant increase at lower redshifts, indicating its dominance in the current epoch of the Universe. The blue line for matter density decreases as redshift decreases, reflecting the transition from a matter-dominated Universe at higher redshifts to a dark energy-dominated Universe at lower redshifts. The cyan line for radiation density is notably higher at early times (high redshifts) and diminishes rapidly as the Universe expands, consistent with the radiation-dominated era in the early Universe. This plot effectively captures the dynamic evolution of the energy components of the Universe, illustrating the transitions from radiation dominance to matter dominance, and finally to dark energy dominance, providing valuable insights into the evolution of the Universe.
V Summary and Conclusion
@@
In this paper, we have delved into the cosmological implications of a modified $f(Q,B)$ gravity model, which integrates both the nonmetricity scalar $Q$ and the boundary term $B$. Our approach adopted the coincident gauge, where the general affine connection vanishes, meaning the covariant derivative reduces to the partial derivative. We then applied Bayesian statistical analysis using MCMC techniques to constrain the model parameters. The analysis was grounded in observational data from CC measurement, the extended Pantheon+ dataset, and BAO measurements. Our results elucidate a smooth transition from a deceleration phase to an accelerating expansion phase in the evolution of the Universe. This transition is critical in understanding the dynamics of cosmic expansion and the role of DE. We developed a numerical approach to predict the redshift behavior of the Hubble expansion rate. This approach was instrumental in constraining the model parameters and understanding the kinematic evolution of the Universe. The $f(Q,B)$ model has been compared with the standard $\Lambda$CDM model, demonstrating its potential as a viable alternative cosmological framework. While the $\Lambda$CDM model has been the cornerstone of modern cosmology, our findings suggest that the $f(Q,B)$ model can replicate the low-redshift behavior of the $\Lambda$CDM model and exhibits notable differences at high redshifts. Our findings align strongly with current cosmological observations of a late-time Universe dominated by DE and undergoing accelerated expansion. This supports the validity of the $f(Q,B)$ model as an alternative explanation for the observed acceleration of the Universe.
@@
A dynamical system analysis framework has been introduced to assess the stability of the model. The identification of a stable critical point using CMT underscores the robustness of the $f(Q,B)$ model. A significant finding of our study is the identification of a stable critical point within the dynamical system of the model, corresponding to the de Sitter phase. The stability of this critical point implies that, given specific initial conditions, the Universe will inherently move towards and stay within the de Sitter phase. This observation aligns with current data indicating a Universe dominated by DE and undergoing late-time accelerated expansion. Future research could delve deeper into the specific initial conditions leading to the de Sitter phase and investigate the influence of the boundary term $B$ on the dynamics of the system. The density parameter plot depicts the transition of the Universe from radiation dominance to matter dominance, and ultimately to dark energy dominance. This offers valuable insights into the evolutionary dynamics of the Universe. Additionally, exploring the implications of this stable critical point for physical quantities like the Hubble parameter would offer valuable insights into the evolution of the Universe. In summary, the $f(Q,B)$ gravity model not only aligns well with current cosmological observations but also provides a comprehensive framework for understanding the accelerated expansion of the Universe. The ability of the model to capture the transition from deceleration to acceleration, identify a stable critical point, and offer a viable alternative to the $\Lambda$CDM model makes it a promising candidate for further exploration in cosmological studies. Our study emphasizes the crucial role of modified gravity theories in understanding the expansion of the Universe and the nature of DE.

### v5

@@ 4 Dynamical System Analysis
$\frac{dY}{dN}$ $=$ $2-3V+\frac{2Z}{3}-\frac{2Zf_{Q}}{3}+3X-2f_{Q}+W-\frac{2f_{Q}^{\prime}}{3}+XZ-YZ\,,$ (49b)
@@
$\frac{dV}{dN}$ $=$ $-4V-2ZV\,,$ (49d)
where $\lambda=\frac{\ddot{H}}{H^{3}}$, we will concentrate on the scenario where $f(Q,B)=f_{0}Q^{m}B^{n}$. The model for this scenario can be expressed using the dynamical variables
@@
and we get the following dependency relation:
$W$ $=$ $-\frac{X}{n}(Z+3)\,,$ (51)
$\Omega_{\text{m}}$ $=$ $-V-2f_{Q}+W+3X+XZ-Y\,,$ (52)
@@
It is possible to eliminate the equations for $W$, $\Omega_{\text{m}}$ and $\Omega_{\text{DE}}$ from our autonomous system using the relations mentioned and Equation (46), resulting in a set of only four equations:
@@
$\frac{dV}{dN}$ $=$ $-4V-2ZV\,.$ (55d)
It is important to note that a dynamical system has a critical point, and this point must be taken into account when analyzing the system,
@@
for the existence condition $m=1-n$. We will now analyze the range of values for $n$, which will result in a stable critical point. While we will not explicitly mention the area of instability, whether it is saddle-like or repulsor-like, it is important to note that the critical point is a de Sitter acceleration phase; therefore, any kind of instability of the critical point is not supported by observations.
@@
$\left\{0,-4,\frac{-n-\sqrt{-2n^{2}-3n}}{n},\frac{-n+\sqrt{-2n^{2}-3n}}{n}\right\}\,.$ (57)
The phase portrait in Figure 6 shows the behavior of a dynamical system near a stable critical point, $\mathcal{P}_{\star}$. As time progresses, trajectories in the phase space tend to move toward $\mathcal{P}_{\star}$, indicating that it is an attractor for the system. This convergence from various initial conditions signifies that small perturbations decay over time, returning the system to the stable state at $\mathcal{P}_{\star}$. The stability of $\mathcal{P}_{\star}$ can be analyzed using the Jacobian matrix evaluated at $\mathcal{P}_{\star}$. The stability of $\mathcal{P}_{\star}$ is ensured when all eigenvalues of the Jacobian possess negative real parts, which aligns with the observed behavior in the phase portrait. The overall dynamics of the system are governed by the differential equations defining it, and the phase portrait provides a graphical representation to visualize these dynamics and understand the long-term behavior of the system. The zero eigenvalues suggest the presence of a center manifold, which requires further analysis to determine overall stability.
4.1 Stability analysis for $\mathcal{P}_{\star}$ by using CMT
CMT is a crucial aspect of dynamical systems theory, focusing on the behavior of systems near fixed points. As detailed by Perko (2013), CMT provides a framework for understanding the stability of these points, especially when traditional linear stability theory falls short due to the presence of zero eigenvalues. This theory reduces the dimensionality of the system near critical points, allowing for a more manageable analysis of stability. When a system approaches a critical point, it tends to behave according to an invariant local center manifold, denoted as $W^{c}$. This manifold is associated with eigenvalues that have zero real parts, capturing the essential dynamics of the system near equilibrium.
Consider a function $f$ belonging to $C^{r}(E)$, where $E$ is an open subset of $\mathbb{R}^{n}$ that includes the origin, and $r\geq 1$. Assume $f(0)=0$ and that the derivative $Df(0)$ possesses $c$ eigenvalues with zero real parts and $s$ eigenvalues with negative real parts, where $c+s=n$. Generally, the system can be reformulated as follows:
@@ 4.1 Stability analysis for $\mathcal{P}_{\star}$ by using CMT
$\dot{y}=By+G(x,y)\,,$
where $A$ is a square matrix with $c$ eigenvalues having zero real parts and $B$ is a square matrix with $s$ eigenvalues having negative real parts and $(x,y)\in\mathbb{R}^{c}\times\mathbb{R}^{s}$. The functions $F$ and $G$ satisfy $F(0)=G(0)=0$ and their derivatives at zero are also zero. Furthermore, there is a small positive value $\epsilon>0$ and a function $g(x)$ in $C^{r}(N_{\epsilon}(0))$, which defines the local center manifold and satisfies certain conditions,
@@
for $|x|<\epsilon$. The center manifold can be derived using the system of differential equations,
@@
for all $x\in\mathbb{R}^{c}$ with $|x|<\epsilon$.
At the critical point $\mathcal{P}_{\star}$, the Jacobian matrix for the autonomous system represented by Equation (55) is as follows:
$J\left(\mathcal{P}_{\star}\right)=\left(\begin{array}[]{cccc}0&1&0&0\\ -3+\frac{3}{n}&-2&-\frac{4}{3}&-3\\ 0&\frac{9}{2n}&0&0\\ 0&0&0&-4\\ \end{array}\right)$ (61)
The eigenvalues of the Jacobian matrix, as presented in Equation (57), are $\lambda_{1}=0$, $\lambda_{2}=-4$, $\lambda_{3}=\frac{-n-\sqrt{-2n^{2}-3n}}{n}$, and $\lambda_{4}=\frac{-n+\sqrt{-2n^{2}-3n}}{n}$. The corresponding eigenvectors are $[\frac{4n}{9-9n},0,1,0]^{T}$, $[-\frac{3n}{11n+3},\frac{12n}{11n+3},-\frac{27}{22n+6},1]^{T}$, $[\frac{2n}{9},\frac{2}{9}\left(\sqrt{-n(2n+3)}-n\right),1,0]^{T}$ and $[\frac{2n}{9},\frac{1}{9}(-2)\left(n+\sqrt{-n(2n+3)}\right),1,0]^{T}$.
Using the CMT, we examine the stability of the critical point $\mathcal{P}_{\star}$. By applying the transformation $X=x-\frac{2n}{3(n-1)}$, $Y=y$, $Z=z$, and $V=v$, we shift this critical point to the origin. The resulting equations in the new coordinate system are as follows:
@@
Upon examining the diagonal matrix in relation to the standard form (Equation (58)), it is clear that the variables $y$, $z$ and $v$ remain stable, while $x$ acts as the central variable. At this critical point, matrices $A$ and $B$ take on the following form:
@@
In the context of CMT, the manifold is characterized by a continuous differential function. Assuming specific functions for the stable variables $y=g_{1}(x)$, $z=g_{2}(x)$, and $v=g_{3}(x)$, we derived the zeroth approximation of the manifold functions using Equation (59)
@@
where the $\mathcal{O}^{2}$ term encompasses all terms that are proportional to the square or higher powers. The following expression gives the center manifold in this scenario:
@@
According to the CMT, the critical point $\mathcal{P}_{\star}$ exhibits stable behavior for $(n<0)\vee(n>1)$.
Figure 6: This graph shows the behavior of a four-dimensional system (Equation (55)) simplified to two dimensions. The parameters are set as $Z=0$, $V=0$, and $n=3.012$.
Figure 6 reveals that a specific point in this two-dimensional representation (the critical point $\mathcal{P}_{\star}$) attracts other points over time, suggesting its stable and attractive nature within the full four-dimensional system. Figure 6 shows the fascinating world of sink trajectories within a dynamical system, visualized through a phase portrait. This point signifies a location where trajectories tend to sink or converge. The critical point $\mathcal{P}_{\star}$ is nonhyperbolic due to the presence of zero eigenvalues and can describe the acceleration of the Universe. It is also an attractor solution, stable in the regions $(n<0)\vee(n>1)$ as determined by CMT. The density parameters for radiation, matter, and DE are $\Omega_{\text{r}}=0$, $\Omega_{\text{m}}=2-\frac{4}{n-1}$, and $\Omega_{\text{DE}}=-1+\frac{4}{n-1}$, respectively, satisfying the constraint in Equation (46). This scenario corresponds to a deceleration parameter $q=-1$ and a total EoS $\omega_{\text{tot}}=-1-\frac{2\dot{H}}{3H^{2}}=-1$, indicating a de Sitter phase and, consequently, an accelerating expansion of the Universe.
@@
Figure 7 shows the evolution of density parameters for DE, matter, and radiation as a function of redshift. The model parameters $m=-2.011$ and $n=3.012$, which are obtained from the parametrization method using MCMC analysis for the CC + Pantheon+ + BAO data sets, are used.
@@
  • The radiation density parameter ($\Omega_{\text{r}}$) is represented by the cyan line, which remains almost constant at zero for small redshift values. This emphasizes its minimal contribution in the present epoch.
@@
In Figure 7, we present the evolution of the density parameters for DE, matter, and radiation as a function of redshift ($z$), utilizing a model parameter value of $m=-2.011$ and $n=3.012$ obtained from our MCMC analysis. The initial conditions for this plot are $X=10^{11}$, $Y=3.29\times 10^{14}$, $Z=0.008$, and $V=4.54\times 10^{-5}$. The magenta line representing DE exhibits a significant increase at lower redshifts, indicating its dominance in the current epoch of the Universe. The blue line for matter density decreases as redshift decreases, reflecting the transition from a matter-dominated Universe at higher redshifts to a DE-dominated Universe at lower redshifts. The cyan line for radiation density is notably higher at early times (high redshifts) and diminishes rapidly as the Universe expands, consistent with the radiation-dominated era in the early Universe. This plot effectively captures the dynamic evolution of the energy components of the Universe, illustrating the transitions from radiation dominance to matter dominance and finally to DE dominance, providing valuable insights into the evolution of the Universe.
5 Summary and Conclusion
@@
In this paper, we have delved into the cosmological implications of a modified $f(Q,B)$ gravity model, which integrates both the nonmetricity scalar $Q$ and the boundary term $B$. Our approach adopted the coincident gauge, where the general affine connection vanishes, meaning the covariant derivative reduces to the partial derivative. We then applied Bayesian statistical analysis using MCMC techniques to constrain the model parameters. The analysis was grounded in observational data from CC measurement, the extended Pantheon+ data set, and BAO measurements. Our results elucidate a smooth transition from a deceleration phase to an accelerating expansion phase in the evolution of the Universe. This transition is critical in understanding the dynamics of cosmic expansion and the role of DE. We developed a numerical approach to predict the redshift behavior of the Hubble expansion rate. This approach was instrumental in constraining the model parameters and understanding the kinematic evolution of the Universe. The $f(Q,B)$ model has been compared with the standard $\Lambda$CDM model, demonstrating its potential as a viable alternative cosmological framework. While the $\Lambda$CDM model has been the cornerstone of modern cosmology, our findings suggest that the $f(Q,B)$ model can replicate the low-redshift behavior of the $\Lambda$CDM model and exhibits notable differences at high redshifts. Our findings align strongly with current cosmological observations of a late-time Universe dominated by DE and undergoing accelerated expansion. This supports the validity of the $f(Q,B)$ model as an alternative explanation for the observed acceleration of the Universe. Our study reveals a significant divergence in the $H_{0}$ measurements derived from the CC + Pantheon+ and CC + Pantheon+ + BAO data sets, as illustrated in Figure 4. This discrepancy underscores the persistent tension surrounding the Hubble constant within the field of cosmology. Resolving these discrepancies is critical for improving our comprehension of the expansion rate of the Universe.
@@
A dynamical system analysis framework has been introduced to assess the stability of the model. The identification of a stable critical point using CMT underscores the robustness of the $f(Q,B)$ model. A significant finding of our study is the identification of a stable critical point within the dynamical system of the model, corresponding to the de Sitter phase. The stability of this critical point implies that, given specific initial conditions, the Universe will inherently move toward and stay within the de Sitter phase. This observation aligns with current data indicating a Universe dominated by DE and undergoing late-time accelerated expansion. Future research could delve deeper into the specific initial conditions leading to the de Sitter phase and investigate the influence of the boundary term $B$ on the dynamics of the system. The density parameter plot depicts the transition of the Universe from radiation dominance to matter dominance and ultimately to DE dominance. This offers valuable insights into the evolutionary dynamics of the Universe. Additionally, exploring the implications of this stable critical point for physical quantities like the Hubble parameter would offer valuable insights into the evolution of the Universe. In summary, the $f(Q,B)$ gravity model not only aligns well with current cosmological observations but also provides a comprehensive framework for understanding the accelerated expansion of the Universe. The ability of the model to capture the transition from deceleration to acceleration, identify a stable critical point, and offer a viable alternative to the $\Lambda$CDM model makes it a promising candidate for further exploration in cosmological studies. Our study emphasizes the crucial role of modified gravity theories in understanding the expansion of the Universe and the nature of DE.